Hello out there to everyone,I need some help with some problems,the first one is,Solve:2n(n+5)=0.The next one is,Find the amount of money in a bank account if the initial deposit is $4,000 with an annual rate of 2.5% for 5 years.The third one is,Use the quadratic formula to solve:5x^2+12=-6x.And the last one is,Use the quadratic formula to solve:-x^2+7x negative10 =0.Thanks to whoever can help.
Answer by stanbon(75887)

the first one is,Solve:2n(n+5)=0
Since the product is zero one of those factors must be zero
It's not the 2.
Either n=0 or n+5=0
So, n=0 of n=-5
------------------------
The next one is,Find the amount of money in a bank account if the initial deposit is $4,000 with an annual rate of 2.5% for 5 years.
After one year you have 4000(1+0.025)
After two years you have 4000(1+0,025)^2
...
After 5 years you have 4000(1+0.025)^5 = 4000*1.025^5
=4000*1.131408...
=$4525.63
-------------------
The third one is,Use the quadratic formula to solve:5x^2+12=-6x
5x^2+6x+12=0
x = [-6+-sqrt(6^2-4*5*12)]/10
x = [-6+-sqrt(-204)]/10
Comment: This would give you complex number solutions and I
have a feeling that you have not studied complex numbers.
Did you post this prolem properly.
------------------------------
And the last one is,Use the quadratic formula to solve:-x^2+7x-10 =0.
x = [-7+-sqrt(7^2-4*-1*-10)]/(-2)
x = [-7+-sqrt(9)]/(-2)
x = [-7+-3]/(-2)
x = [-7+3]/(-2) or x = [-7-3]/(-2)
x = 2 or x = 5
Cheers,
Stan H.
---------------